OpenStudy (anonymous):

a fair die is tossed five times. what is the probability of tossing a 6 exactly four times?

OpenStudy (fibonaccichick666):

what is the probability of one time?

OpenStudy (anonymous):

It is a 1/6 chance each time, so to roll that 4 times in a row is (1/6)^4

OpenStudy (anonymous):

Woops, I think I just ruined Fibonacci's strategy x.x

OpenStudy (fibonaccichick666):

it's ok

OpenStudy (anonymous):

I'm abit confused okay so we know that there's only one 6 on a die

OpenStudy (anonymous):

giving you a chance of 1/6 on every roll

OpenStudy (fibonaccichick666):

correct

OpenStudy (fibonaccichick666):

now let's figure out the probability for getting two 6's in a row

OpenStudy (fibonaccichick666):

so if there is a \(\frac{1}{6}\) chance the first time what is the chance of getting a six on the second roll?

OpenStudy (anonymous):

1/5?

OpenStudy (anonymous):

wait srry

OpenStudy (anonymous):

There are many possibilities, it can be one of 6 things after you rolled the first 6.

OpenStudy (anonymous):

And the first 6 is one of six possibilities as well, and all of the other possibilities have those same secondary possibilities.

OpenStudy (anonymous):

If you want to attack thus part by brute force, think of it like this: 1st roll 2nd roll 1 1 2 2 3 3 4 4 5 5 6 6 Each of the first roll's results can be linked to each of the second's, right?

OpenStudy (anonymous):

So how many possible outcomes can you have?

OpenStudy (anonymous):

36

OpenStudy (anonymous):

or wait 6^5?

OpenStudy (anonymous):

No you had it right the first time, 36 possibilities. And how many of those includes a 6 and a 6?

OpenStudy (anonymous):

There is only one instance where you can have a 6 and a 6.

OpenStudy (anonymous):

So that chance would be...?

OpenStudy (anonymous):

1/6

OpenStudy (anonymous):

Think about the whole situation. 36 outcomes, only 1 of them you want. What is the chance you'll get that outcome?

OpenStudy (anonymous):

I'm so stupid at probabilities, forgive me :/

OpenStudy (anonymous):

I need to go like really soon, so I'll tell you that this part is (1 outcome)/(36 outcomes), or 1/36

OpenStudy (anonymous):

ohhh wow I get it

OpenStudy (anonymous):

Luis, now what are you doing? I don't understand where you got any of that -.-

OpenStudy (anonymous):

I don't think Luis is correct here, just listen to what I said. And Fibonacci can you finish this one? I really need to go.

OpenStudy (anonymous):

thanks for the help once again, :)

OpenStudy (anonymous):

I kind of get it

OpenStudy (rajee_sam):

since it is rolled 5 times wont you consider the probability of not getting a 6 the last time which is 5/6 and multiply that to 1/6^4. so the answer is 5/6^5??

OpenStudy (anonymous):

So did you figure it out yet?

OpenStudy (zarkon):

this is binomial... \[{5\choose 4}\left(\frac{1}{6}\right)^{4}\left(\frac{5}{6}\right)^1\]

OpenStudy (anonymous):

uhh, yes, @smokeydabear , I do know what I'm talking about. What i wrote is just the expression for each toss so he would finish the rest. But im going to keep this short cuz its way too long. You must distinguish between tosses and outcomes of each toss. But remember that when asked for the probability of exactly "something" you have to multiply the possibility that the "something " happens times the possibility that it does not happen. Think of it like the outcomes of tossing a coins. But this is only for one toss .But since there are 5 tosses, multiply this by 5. To sum it all up , the answer is

OpenStudy (anonymous):